The scores on a national achievement exam are normally distributed with a mean of 600 and a standard deviation of 120. If a student is selected at random, find the probability that the student scored above 660


1
Expert's answer
2021-03-26T04:29:46-0400



Since "\mu=600" and "\sigma =120" we have:

"P (X>660) = P(X-\mu>660-600) = P(\dfrac{X-\mu}{\sigma}-\dfrac{660-600}{120})"

Since "\dfrac{X-\mu}{\sigma}" and "\dfrac{660-600}{120}" = 0.5 we have:

P(X > 660) = P(Z > 0.5)

Use the standard normal table to conclude that:

P(Z > 0.5) = 0.3085